RC meeting and Run Plan, Friday February 10th

From SBS wiki
Jump to: navigation, search

Return to the RC Meeting Calendar

RC Report


  • Ran with beam most of yesterday, collecting 1.2C
  • Beam tripping increased and MCC had to drop us to 30 uA during owl.
  • This morning MCC did some beam studies and feels confident that the trips will be reduced now.
  • Provakar and Arun investigated the BBCal FIFO but discovered that the upstream summing module is the actual issue. This will take a 1 hour access to fix.


Current Status

  • Running 45 uA on He3

Run Plan

  • Continue running 45 uA on He3

Data Collection and Polarization Status

  • 25.2C of uncorrected charge for this kinematic (27.9C of corrected charge)
  • Last NMR 52.7%

Other Topics/Status Updates

  • Anything else?

Requested Accesses to the Hall

  • Arun or Provakar to replace the BBCal summing module
  • Jack would like to collect some cables
  • GEM team need like a 3-4hour access for the SBS GEMs DAQ troubleshooting